PRINTER VERSION BACK NEXT Problem 2.23 A soccer ball of mass 0.39 kg is rolling with velocity <0, 0, 1.5> m/s, when you kick it. Your kick delivers an impulse of magnitude 1.6 N.s in the -x direction. The net force on the rolling ball, due to the air and the grass, is 0.27 N in the direction opposite to the direction of the balls momentum. Using a time step of 0.5 s, find the position of the ball at a time 1.5 s after you kick it, assuming that the ball is at the origin at the moment it is kicked. Use the approximation v avg ~ะ P/m. (Express your answer in vector form.) > m LINK TO TEXT By accessing this Question Assistance, you will learn while you earn points based on the Point Potential Policy set by your instructor. Question Attempts: 0 of 10 used SAVE FOR LATER SUBMIT ANSWER Earn Maximum Points available only if you answer this question correctly in three attempts or less
Answer #1

Using impulse momentum

I = m(v-u)

I = impulse, m , mass, v and u = final and initial velocity

-1.6i = 0.39(v - 1.5k)

v = -4.1i + 1.5k m/s

This v will become u for next part

The unit vector of v will give direction of ball and help finding acceleration

v^ = 1/(4.1^2 + 1.5^2)^0.5 x (-4.1i + 1.5k)

= -0.94i + 0.34k

The direction opposite to that will be direction of acceleration

a^= 0.94i - 0.34k

a = F/m x a^

= 0.27/0.39 x (0.94i-0.34k)

= 0.65i - 0.24k m/s^2

Using v = u + at

V = (-4.1+1.5) + (0.65i-0.24k)x1.5

= -3.125i + 1.14k m/s